A satellite dish in the shape of a paraboloid is 10 ft across, and 4 ft deep at its vertex. How far is the receiver from the vertex, if it is placed at the focus? Round off your answer to 2 decimal places.
1
Expert's answer
2020-10-19T16:45:19-0400

let c the distance between the vertex and the focus.

the equation for a vertical parabola:

"y = \frac{1}{4c}(x-h)^2 + k"

When the center is at origin

then this equation becames

"y = \frac{1}{4c}x^2"

.

The problem gives you a point on the parabola: (10,4)

put it in and solve for 'c':

"4=\frac{1}{4c}10^2"


"c=\frac{10^2}{16}"


c "=\frac{100}{16}=6.25 feet"

So the required answer is 6.25 feet